The company wants to sell t-shirts and owner  wants to be 95% certain of the results, and she wants to be within $1,000 of the true mean. When she was researching online, she found that the standard deviation of the mean t-shirt sales for businesses similar to hers was $2,200.

How many should she get data from?
Answer #1

Solution :

Given that,

standard deviation = =  \sigma =2200

Margin of error = E = 1000

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.96

sample size = n = [Z\alpha/2* \sigma / E] 2

n = ( 1.96* 2200 / 1000 )2

n =18.59